OpenStudy (anonymous):

If fog (x) = 2x^2 +6x+5 and f(x) = 2x^2 -2x+1, what is g(x)? (Hint, use the quadratic formula)

OpenStudy (anonymous):

\[g=g(x)\]\[fog(x)=f(g(x))=f(g)=2g^2-2g+1=2x^2+6x+5\]\[g^2-g=x^2+3x+2\]does that help?

OpenStudy (anonymous):

oh! Makes sense. Thank you! :)